OpenStudy (narusamisty):

Please help :) Will give medal and fan :D Use technology or a z-distribution table to find the indicated area. The lengths of green beans for sale at a supermarket are normally distributed with a mean of 11.2 cm and a standard deviation of 2.1 cm. Consider a bag of 150 green beans. How many green beans will be 13 cm or shorter?

OpenStudy (jjjjjjjjjjjjhujljnjb):

he Rong he bad anser is 4

OpenStudy (narusamisty):

?

OpenStudy (agent0smith):

Find the z score first, from (score - mean)/SD: \[\large z = \frac{ x - \mu }{ \sigma }\] Then use a z table to find the probability. Then you can multiply by 150 beans.

OpenStudy (narusamisty):

I multiplied 0.8051*150 and I got 120.765

OpenStudy (narusamisty):

I got 121

OpenStudy (narusamisty):

is that correct? :)

OpenStudy (narusamisty):

Could you help me with this one too :) Use the probability distribution graph to answer the question. P(X≤a)=0.75P(X≤a)=0.75 What is the value of a?

OpenStudy (narusamisty):

*P(X≤a)=0.75

OpenStudy (narusamisty):

@agent0smith

OpenStudy (narusamisty):

i think a=1? :)

OpenStudy (agent0smith):

That one is really easy. The area of the rectangle for P(X≤a) must be 0.75. You have the height... it's 0.25. So work out what a is.

OpenStudy (narusamisty):

alright thankyou!!!!

OpenStudy (agent0smith):

a=1 is not correct btw.

OpenStudy (narusamisty):

ah ok i will make sure i fix that thanks :)

OpenStudy (agent0smith):

It's easy, work out what a is... the rectangle area must be 0.75, and the height is 0.25, so what x value gives an area of 0.75?

OpenStudy (narusamisty):

5?

OpenStudy (agent0smith):

Nope... look at the graph. Tell me the area between x=1 and x=2.

OpenStudy (narusamisty):

would it be 0.25 sorry if im not getting this right I've just started learning this :(

OpenStudy (agent0smith):

Yes. But this is basic geometry, don't make excuses :P If the area between x=1 and 2 is 0.75, what about the area between x=1 and x=3? x=1 and x=4 x=1 and x=5?

OpenStudy (narusamisty):

1.5 2 2.25 3 is it wrong that i multiplied 0.75 by the spaces between the number?

OpenStudy (agent0smith):

Why would you multiply 0.75 by anything? x=1 and x=3? - what is the base of the rectangle? what is the height?

OpenStudy (agent0smith):

If one 'column' is 0.25 (x=1 to x=2), then what is the area of two 'columns' (x=1 to x=3)?

OpenStudy (narusamisty):

0.5?

OpenStudy (agent0smith):

You should know for sure. This is easy. It's area of rectangles!

OpenStudy (narusamisty):

A=w*l know the length is 0.25 but like i said i dont know what to put in for the width

OpenStudy (narusamisty):

what i did before is that i counted the columns and the i multiplied it by 0.25 thats how i the got 0.5 :(

OpenStudy (agent0smith):

You can work out the width... with subtraction :P what is the width if you're looking between x=1 and x=2? It's 1. What's the width between x=1 and x=3?

OpenStudy (narusamisty):

2

OpenStudy (agent0smith):

Yes. Now use that method to find the value of a: P(X≤a)=0.75

OpenStudy (narusamisty):

the width between 1 and 5 is 4 and i multiplied that with the length 0.25 but i end up getting 1 :/ l*w=A 0.25*4=1

OpenStudy (narusamisty):

Ah ok i understand it now i saw what i did wrong

OpenStudy (agent0smith):

P(X≤a)=0.75 So then the value of a is...?

OpenStudy (narusamisty):

3 right you asked me what value gave me 0.75

OpenStudy (agent0smith):

Not quite. a is the x value. What x value gives an area of 0.75.

OpenStudy (narusamisty):

so 3 isn't the x value? because i used the formula and I multiplied 0.25 by 3 (the x value )which would then give me 0.75 :/

OpenStudy (agent0smith):

P(X≤a)=0.75 http://assets.openstudy.com/updates/attachments/57e0bbcee4b0860740296b37-narusamisty-1474349449600-capture.png Look at the graph. Which x value gives an area 0.75?

OpenStudy (narusamisty):

oh its 4... ._.

OpenStudy (agent0smith):

Yes.

OpenStudy (narusamisty):

thankyou for helping me. :)

OpenStudy (agent0smith):

Welcome :)